# Diagnosing Quantum Many-body Chaos in Non-Hermitian Quantum Spin Chain via Krylov Complexity

### v1

@@ Abstract
We investigate the phase transitions from chaotic to non-chaotic dynamics in a quantum spin chain with a local non-Hermitian disorder, which can be realized with a Rydberg atom array setting. As the disorder strength increases, the emergence of non-chaotic dynamics is qualitatively captured through the suppressed growth of Krylov complexity, and quantitatively identified through the reciprocity breaking of Krylov space. We further find that the localization in Krylov space generates another transition in the weak disorder regime, suggesting a weak ergodicity breaking. Our results closely align with conventional methods, such as the entanglement entropy and complex level spacing statistics, and pave the way to explore non-Hermitian phase transitions using Krylov complexity and associated metrics.
@@
Introduction.— Diagnosis of chaos emerging in quantum systems attracts a growing interest, with a variety of measures including Lyapunov exponents [1, 2, 3], out-of-time-order correlations [4, 5, 6], Loschmidt echo [7, 8, 9], as well as spectral characteristics revealed through level spacing distributions [10, 11, 12, 13, 14] and spectral form factors [15, 16, 17]. Recently, Krylov complexity has emerged as an innovative approach to the study of quantum chaos [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33]. Krylov complexity quantifies the spreading of quantum operators [18] or states  [20] in Krylov space, capturing key features of the underlying dynamics.
@@
Figure 1: (a) Disordered non-Hermitian XY model, with XY coupling $J$, transverse field $h$, site-dependent detuning $\Delta_{j}\in[-W_{\Delta},W_{\Delta}]$, and dissipation $\gamma_{j}\in[-W_{\gamma},W_{\gamma}]$. (b) Tunneling rates $b_{j}$ and $c_{j}$ are not conjugate for non-Hermitian systems. (c) Sketch of the phase diagram with increasing $W_{\gamma}$. (d) Krylov wave functions in chaotic regimes with delocalized (left) and localized (middle) profiles, and non-chaotic (right) regimes, at mid-term $t=300$ (solid blue) and long-term $t=1000$ (dashed red). See text for details.
@@
A pivotal insight in this field is the operator growth hypothesis [18], which suggests that information scrambling and chaos can be identified by observing exponential growth in Krylov complexity at early times. The exponential growth is attributed to the linearity of Lanczos coefficients, which govern the tunneling rates in Krylov space. In contrast, another hypothesis focuses on the long-term dynamics and proposes a Tsallis $q$-log behavior of the Lanczos coefficients throughout the entire Krylov space [24, 32], where chaos can be signified by the linear growth in complexity at intermediate times, followed by a peak before saturation [20, 24, 27]. Additionally, the fluctuation in Lanczos coefficients, known as Krylov variance, has been shown to correlate with the delocalization of the Krylov wave function and, consequently, with ergodic behaviors [19, 21, 25, 29, 28, 33]. Despite extensive investigations in the Hermitian regime, challenges emerge when extending to open quantum systems. While Krylov complexity appears to be significantly reduced under dissipation [34, 35, 36, 37, 38, 39], the Lanczos coefficients deviate from the established hypotheses for Hermitian systems, particularly when the absence of Hermiticity of Krylov space is overlooked. Furthermore, the critical behavior and universal properties near phase transitions in open systems remain largely unexplored.
@@
In this work, we investigate Krylov complexity of a non-Hermitian XY model with random local gain and loss. We show that the evolution of Krylov wave functions and complexity growth are indicative of phase transitions of quantum chaos. Through a closer examination of Lanczos coefficients, we uncover that the Krylov variance elucidates critical points and universal features for localization of Krylov wave function which indicates a weak ergodicity breaking. We introduce a metric of reciprocity to underscore the suppression of quantum chaos, where the onset of non-reciprocity marks the corresponding critical point. The critical points determined using our method are quantitatively consistent with established metrics, including entanglement entropy variance [40, 41] and complex level spacing statistics [12, 13, 14]. Our framework is also applicable to other ergodicity-breaking phenomena, including many-body localization [21, 42], quantum scars [43], and Hilbert space fragmentation [44], and holds potential for extensions into machine learning applications [45]. The non-Hermitian XY model can be realized with Rydberg atom array quantum simulators. Our work thus opens avenues to experimentally explore complexities with observables such as Loschmidt echo [7, 46].
@@
Model.— We study a one-dimensional disordered XY spin chain under both transverse and longitudinal fields, as depicted in Fig. 1(a). The Hamiltonian is given by
@@
where $\hat{\sigma}_{j}^{x}$, $\hat{\sigma}_{j}^{y}$, $\hat{\sigma}_{j}^{z}$ are Pauli matrices acting on the $j$-th site. The longitudinal field is random and complex, $\mathcal{D}_{j}=\Delta_{j}+{\rm i}\gamma_{j}$, where the real part $\Delta_{j}$ and imaginary part $\gamma_{j}$ give the detuning, and the gain or loss, respectively. The parameters $\Delta_{j}$ and $\gamma_{j}$ are independently and uniformly distributed within $[-W_{\Delta},W_{\Delta}]$ and $[-W_{\gamma},W_{\gamma}]$, correspondingly. In the following calculations, we scale the Hamiltonian with the XY interaction strength $J$. To explore the non-Hermitian effects, we vary $W_{\gamma}$ while fixing $h=0.5$ and $W_{\Delta}=1$ throughout the work. The model can be realized with Rydberg atom arrays, which will be discussed later.
@@
In the Krylov space, the dynamics are mapped to a single-particle tight-binding model [see Fig. 1(b)]. The non-Hermitian Hamiltonian displays vastly different dynamics as the disorder strength $W_{\gamma}$ increases. Specifically, two phase transitions are observed, where a suppression of spreading within finite Krylov subspace takes place at $W_{\gamma}=W_{L}\approx 0.0351$ revealing a localization feature, and a transition from chaotic to non-chaotic dynamics at $W_{\gamma}=W_{C}\approx 1.647$. The phase diagram is depicted in Fig. 1(c). These phases can be captured by the spreading of the Krylov wave function, illustrated in Fig. 1(d) [see Supplemental Materials (SM) for more examples]. For $W_{\gamma}<W_{L}$, the Krylov wave function is fully delocalized across the entire Krylov space at late times ($t=1000$). For $W_{\gamma}>W_{L}$, conversely, a weak breaking of ergodicity may be indicated by the localized Krylov wave function with exponentially small occupation for $n\gtrsim\mathcal{O}(L^{2})$. In the non-chaotic regime, $W_{\gamma}<W_{C}$, the wave function is further confined below $n\lesssim\mathcal{O}(L)$. We also observe prethermal behavior in the Krylov delocalized regime, where the wave function is confined near $n\approx 0$ at mid-term times ($t=300$), contrasting with other phases. Krylov complexity and Lanczos coefficients capture critical behaviors around the phase transition points. This allows us to determine the critical points through a number of measures, which will be discussed in detail in the following.
Figure 2: (a) Krylov complexity growth. With $W_{\Delta}=0.5$, $W_{\gamma}$ varies from $0.001$ to $5$ across system sizes (a1) $L=6$, (a2) $L=8$, and (a3) $L=10$. Dashed lines are the quadratic ($\sim t^{2}$) and dash-dot lines the linear (tsimilar-$\sim t$) trends, respectively. (b) Lanczos coefficients for $L=10$ and $W_{\gamma}=$ (b1) 0.0118, (b2) 0.2, (b3) 1.0, and (b4) 3.0. Upper panels show $\langle|b_{n}|\rangle$ (solid blue) and $\langle|c_{n}|\rangle$ (dashed red), with dotted lines representing $\sim n$ for $n<L$, and $\sim\sqrt{1-n/2^{L}}$ for $n>L$. Lower panels show $\langle\cos\theta_{n}\rangle$. (c) The Krylov variance given by Eq. (4), and the inset is the finite-size scaling, from which we obtain $W_{L}=0.0351$. (d) Krylov reciprocity given by Eq. (5). The inset is the finite-size scaling yielding $W_{C}=1.647$. Simulations are sampled and averaged over 1000, 600, and 400 trajectories for $L=6,8$, and $10$, respectively.
@@
Krylov complexity.— The mathematical framework for representing quantum models in Krylov space involves a tridiagonalization process [47]. Krylov space is constructed by iteratively applying the Hamiltonian to an initial state vector and orthogonalizing it against the preceding basis vectors. This process is efficiently implemented using the bi-Lanczos algorithm, which is capable of dealing with non-Hermitian systems [36, 38, 37]. In Krylov space, the evolution of the quantum system is encapsulated as an effective single-particle tight-binding model on a semi-infinite chain [18]. The Krylov wave function $\varphi_{n}$ is governed by the discrete equation (see SM for details),
$i\dot{\varphi}_{n}=b_{n}\varphi_{n-1}+a_{n}\varphi_{n}+c_{n+1}\varphi_{n+1},$ (2)
@@
where $\varphi_{n}$ represents the distribution over the Krylov space. The Lanczos coefficients $b_{n}$ and $c_{n}$ are generally not complex conjugates, due to the non-Hermiticity of the Hamiltonian. The initial state is $\varphi_{0}=1$ and $\varphi_{n}=0$ for $n>0$.
@@
Krylov complexity quantifies the spreading of the Krylov wave function by its center of mass, defined as
@@
The growth of Krylov complexity is a rich, multistage phenomenon. Figures 2(a1-a3) illustrate the complexity growth for system sizes $L$ = 6, 8, and 10 with the initial state $\left|\psi_{0}\right\rangle=\left|+\right\rangle^{\otimes L}$, where $\left|+\right\rangle=(\left|\uparrow\right\rangle+\left|\downarrow\right\rangle)/\sqrt{2}$. In case of Hermitian systems ($W_{\gamma}=0$), the Krylov complexity typically shows generic quadratic-linear-peak-saturate shape, reminiscent of the slope-dip-ramp-plateau profile of the spectral form factor [24]. In our non-Hermitian spin chain model, we find distinctive features beyond these trends, indicating Krylov localization and chaos-non-chaos transition as $W_{\gamma}$ increases.
The short-term growth of the Krylov complexity is universally quadratic, $C_{K}(t)\sim(1+W_{\gamma}^{2})t^{2}$ (see SM). The absence of exponential growth suggests a deviation from the operator growth hypothesis, as also corroborated by the nonlinear behavior of Lanczos coefficients [Fig. 2(b1-b4)]. Thus, it precludes chaos diagnosis using the early time dynamics in our model.
On the intermediate time scale, the Krylov complexity grows nonlinearly, deviating from the linear growth associated with spectral rigidity reported in previous studies [24]. Notably, for $W_{\gamma}<W_{C}$, $C_{K}$ is bounded by a linear envelope, and its long-term saturation value increases as $W_{\gamma}$ decreases. These phenomena indicate that chaos suppression correlates with confined spreading of Krylov wave function [Fig. 1(d)]. In chaotic regimes, the Krylov wave functions extend beyond the system size, as a result of the involvement of many excited states. Ref. [24] suggests a linear ramp of the Krylov complexity due to the Lanczos coefficients following $b_{n}\propto\sqrt{1-n/2^{L}}$, which is a special case of Tsallis $q$-log statistics with $q=0$ [32]. In contrast, we find $q>0$ in our results (see SM), as shown in Figs. 2(b1-b4). The mid-term ramp, on the other hand, indicates spectral rigidity as it is constrained by a linear envelope [Fig. 2(a)].
In the Krylov delocalized regime ($W_{\gamma}<W_{L}$), the Krylov complexity shows prethermal behaviors, where oscillations and a low saturation plateau emerge before a rapid increase to higher values. This behavior stems from dynamics confined to a small Krylov subspace due to small overlaps between the initial state and eigenstates with large imaginary components. For example, for $W_{\gamma}=0.022$ and at $t=300$, the system forms a prethermal state where the Krylov wave function is localized at $n<\mathcal{O}(L)$ [Fig. 1(d)]. At longer times $t\gtrsim 10^{3}$, $C_{K}$ saturates as the Krylov wave function completely delocalizes. As system size increases, $C_{K}$ exhibits damped oscillations, and the saturation time grows significantly.

### v2

@@ Abstract
We investigate the phase transitions from chaotic to nonchaotic dynamics in a quantum spin chain with a local non-Hermitian disorder, which can be realized with a Rydberg atom array setting. As the disorder strength increases, the emergence of nonchaotic dynamics is qualitatively captured through the suppressed growth of Krylov complexity, and quantitatively identified through the reciprocity breaking of Krylov space. We further find that the localization in Krylov space generates another transition in the weak disorder regime, suggesting a weak ergodicity breaking. Our results closely align with conventional methods, such as the entanglement entropy and complex level spacing statistics, and pave the way to explore non-Hermitian phase transitions using Krylov complexity and associated metrics.
I Introduction
@@
Diagnosis of chaos emerging in quantum systems attracts a growing interest, with a variety of measures including Lyapunov exponents [1, 2, 3], out-of-time-order correlations [4, 5, 6], Loschmidt echo [7, 8, 9], and spectral characteristics revealed through level spacing distributions [10, 11, 12, 13, 14, 15, 16] and spectral form factors [17, 18, 19]. Recently, Krylov complexity has emerged as an innovative approach to the study of quantum chaos [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41]. Krylov complexity quantifies the spreading of quantum operators [20] or states [22] in Krylov space, capturing key features of the underlying dynamics.
@@
Figure 1: (a) Disordered non-Hermitian XY model, with XY coupling $J$, transverse field $h$, site-dependent detuning $\Delta_{j}\in[-W_{\Delta},W_{\Delta}]$, and dissipation $\gamma_{j}\in[-W_{\gamma},W_{\gamma}]$. (b) Illustration of the Krylov chain, which maps the physical system to the Krylov space. The Krylov wave function is initially localized at the first site and then propagate throughout the chain. Tunnelling rates $b_{j}$ and $c_{j}$ are generally not conjugate for non-Hermitian systems. (c) Sketch of the phase diagram with increasing $W_{\gamma}$. The dynamics exhibits a chaos to nonchaos phase transition as the disorder term $W_{\gamma}$ increases. At $W_{\gamma}=0$, the system follows Gaussian orthogonal ensemble (GOE) statistics, transitioning to the AI class near $W_{L}$ and then approaching Poisson statistics for very large $W_{\gamma}$. In Krylov space, wave functions remain delocalized for $W_{\gamma}<W_{L}$ at long times, but localized for $W_{\gamma}>W_{L}$. The first few tunneling coefficients, $b_{j}$ and $c_{j}$, show a transition from reciprocal to nonreciprocal regimes at $W_{C}$, which enhances localization. (d) Krylov wave functions in chaotic regimes with delocalized (left) and localized (middle) profiles, and nonchaotic (right) regimes, at mid-term $t=300$ (solid blue) and long-term $t=1000$ (dashed red). See text for details.
@@
A pivotal insight in this field is the operator growth hypothesis [20], which suggests that information scrambling and chaos can be identified by observing exponential growth in Krylov complexity at early times. The exponential growth is attributed to the linearity of Lanczos coefficients, which govern the tunneling rates in Krylov space. In contrast, another hypothesis focuses on the long-term dynamics and proposes a Tsallis $q$-log behavior of the Lanczos coefficients throughout the entire Krylov space [26, 37, 41], where chaos can be signified by the linear growth in complexity at intermediate times, followed by a peak before saturation [22, 26, 32]. Additionally, the fluctuation in Lanczos coefficients, known as Krylov variance, has been shown to correlate with the delocalization of the Krylov wave function and, consequently, with ergodic behaviors [21, 23, 27, 34, 33, 38]. Despite extensive investigations in the Hermitian regime, challenges emerge when extending to open quantum systems. While the complexity appears to be significantly reduced under dissipation [42, 43, 44, 45, 46, 47, 48, 49], the Lanczos coefficients deviate from the established hypotheses for Hermitian systems, particularly when the absence of Hermiticity of Krylov space is overlooked. Furthermore, the critical behavior and universal properties near phase transitions in open systems remain largely unexplored.
@@
In this work, we investigate Krylov complexity of a non-Hermitian XY model with random local gain and loss. We show that the evolution of Krylov wave functions and complexity growth are indicative of phase transitions of quantum chaos. Through a closer examination of Lanczos coefficients, we find that the Krylov variance elucidates critical points and universal features for localization of the wave function. This indicates a weak ergodicity breaking as the quantum state can only evolve within a confined Krylov subspace. Additionally, in the delocalization regime, the Krylov complexity growth behavior exhibits a prethermal plateau before saturation, which links to the Hermiticity breaking [50]. To identify transitions between chaotic and nonchaotic phases, we introduce a metric of reciprocity, where the onset of nonreciprocity of Krylov space marks the corresponding critical point. Our results are consistent with the complex level spacing statistics [12, 13, 14], which reflect the conjectures of level repulsion in quantum chaos by Bohigas, Giannoni, and Schmit for closed systems [51] and Grobe, Haake, and Sommers for open systems [52]. We have further compared the critical points using our method with the variance of entanglement entropy [53, 54], which are also qualitatively consistent.
Our framework is applicable to other ergodicity-breaking phenomena, including many-body localization [23, 55, 56], quantum scars [57, 58], and Hilbert space fragmentation [59, 60], and holds potential for extensions into machine learning applications [61]. The non-Hermitian XY model can be realized with Rydberg atom array quantum simulators [62, 63, 64] and circuit-based quantum simulators [65, 66, 67, 68]. Our work thus opens avenues to experimentally explore complexities with observables such as Loschmidt echo [7, 69, 70, 71, 72].
This paper is organized as follows. In Sec. II, we describe the non-Hermitian XY model and briefly show the phase transitions about quantum chaos. In Sec. III, we present the concept and definition of Krylov complexity and relevant quantities, and show the overall features of complexity growth under different parameters. In Sec. IV, we use the Krylov variance and reciprocity to quantitatively identify the critical behaviors of phase transitions. Finally, in Sec. V, we compare our results with the conventional signatures including the complex level spacing (CSR) ratio and entanglement entropy.
II Model
@@
We study a one-dimensional disordered XY spin chain under both transverse and longitudinal fields, as depicted in Fig. 1(a). The Hamiltonian is given by
@@
where $\hat{\sigma}_{j}^{x}$, $\hat{\sigma}_{j}^{y}$, and $\hat{\sigma}_{j}^{z}$ are Pauli matrices acting on the $j$th site. The longitudinal field is random and complex, $\mathcal{D}_{j}=\Delta_{j}+{\rm i}\gamma_{j}$, where the real part $\Delta_{j}$ and imaginary part $\gamma_{j}$ give the detuning and the gain or loss, respectively. The parameters $\Delta_{j}$ and $\gamma_{j}$ are independently and uniformly distributed within $W_{\Delta},W_{\Delta}]$ and $W_{\gamma},W_{\gamma}]$, correspondingly. In the following calculations, we scale the Hamiltonian with the XY interaction strength $J$. To explore the non-Hermitian effects, we vary $W_{\gamma}$ while fixing $h=0.5$ and $W_{\Delta}=1$ throughout the work. The model can be realized with Rydberg atom arrays, which will be discussed later.
@@
In the Krylov space, the original physical model is mapped to a single particle moving on a semi-infinite chain. Starting from the left end, the particle is hopping on the chain, giving the so-called Krylov wave function. As a result, the complexity growth is depicted by the propagation of Krylov wave function [see Fig. 1(b)]. The basis of Krylov space can be generated by various Krylov algorithms (see Appendix A for details). The non-Hermitian Hamiltonian displays vastly different dynamics as the disorder strength $W_{\gamma}$ increases. Specifically, two phase transitions are observed, where a suppression of spreading within finite Krylov subspace takes place at $W_{\gamma}=W_{L}\approx 0.0351$ revealing a localization feature, and a transition from chaotic to nonchaotic dynamics at $W_{\gamma}=W_{C}\approx 1.647$. The phase diagram is depicted in Fig. 1(c). These phases can be captured by the spreading of the Krylov wave function, $\varphi_{n}$, illustrated in Fig. 1(d) (see Appendix B for more examples). For $W_{\gamma}<W_{L}$, $\varphi_{n}$ is fully delocalized across the entire space at late times ($t=1000$). For $W_{\gamma}>W_{L}$, conversely, a weak breaking of ergodicity may be indicated by localization with exponentially small occupation for $n\gtrsim\mathcal{O}(L^{2})$. In the nonchaotic regime, $W_{\gamma}<W_{C}$, $\varphi_{n}$ is further confined below $n\lesssim\mathcal{O}(L)$. We also observe prethermal behavior in the delocalized regime, where $\varphi_{n}$ is confined near $n\approx 0$ at mid-term times ($t=300$), contrasting with other phases. Krylov complexity and Lanczos coefficients capture critical behaviors around the phase transition points. This allows us to determine the critical points through a number of measures, which will be discussed in detail in the following.
Figure 2: (a) Krylov complexity growth. $W_{\gamma}$ varies from $0.001$ to $5$ for colors from red to blue (from top to bottom in the long-time limit $t\sim 10^{4}$). We have fixed $W_{\Delta}=0.5$ and system size $L=10$. The dashed line is the quadratic ($\sim t^{2}$) trend and dotted line is the linear ($\sim t$) trend. The scaling of the Krylov complexity at (b) $t=300$ and (c) $t=10^{4}$ for different system sizes, and the colors, from red to blue, indicates the values of $W_{\gamma}=(0.005,0.026,0.06,0.1,0.2,0.8,1.8,4.8)$, in accordance with panel (a). (d), (e) The Krylov inverse participation ratios at the same times as panels (b) and (c).
Early-time growth Intermediate growth Saturated value Krylov localization
$W_{\gamma}<W_{L}$ Quadratic Plateau $\to$ Abrupt increase High Very high $\to$ low
$W_{L}<W_{\gamma}<W_{C}$ Quadratic Linear envelope Transitional High
$W_{\gamma}>W_{C}$ Quadratic Rapid saturation Low Very high
Table 1: A summary of major features of the complexity growth in different regimes of $W_{\gamma}$.
III Krylov complexity
@@
The mathematical framework for representing quantum models in Krylov space involves a tridiagonalization process [73]. Krylov space is constructed by iteratively applying the Hamiltonian to an initial state vector and orthogonalizing it against the preceding basis vectors. This process is efficiently implemented using the bi-Lanczos algorithm, which is capable of dealing with non-Hermitian systems [45, 46, 48, 47]. In Krylov space, the evolution of the quantum system is encapsulated as an effective single-particle tight-binding model on a semi-infinite chain [20]. The wave function $\varphi_{n}$ is governed by the discrete equation (see Appendix A for details),
${\rm i}\frac{\partial}{\partial t}\varphi_{n}=b_{n}\varphi_{n-1}+a_{n}\varphi_{n}+c_{n+1}\varphi_{n+1},$ (2)
@@
where $\varphi_{n}$ represents the wave function of the Krylov space. The Lanczos coefficients $b_{n}$ and $c_{n}$ are generally not complex conjugates, due to the non-Hermiticity of the Hamiltonian. The initial values of the Krylov wave function amplitudes are $\varphi_{0}=1$ and $\varphi_{n}=0$ for $n>0$.
@@
Krylov complexity quantifies the spreading of $\varphi_{n}$ by its center of mass, defined as
@@
We note that the norm of the wave function is not naturally conserved, but we always normalize the Krylov wave function $\varphi_{n}$ before calculating the Krylov complexity. The growth of complexity is a rich, multistage phenomenon. Figure 2(a) illustrates the complexity growth for system sizes $L=10$ with the initial state $\left|\psi_{0}\right\rangle=\left|+\right\rangle^{\otimes L}$, where $\left|+\right\rangle=(\left|\uparrow\right\rangle+\left|\downarrow\right\rangle)/\sqrt{2}$. More examples of the complexity growth for different system sizes are shown in Appendix B. In case of Hermitian systems ($W_{\gamma}=0$), the complexity typically shows generic quadratic-linear-peak-saturate shape, reminiscent of the slope-dip-ramp-plateau profile of the spectral form factor [26]. In our non-Hermitian spin chain model, instead of these trends, we find distinctive features, including Krylov localization, and chaos-nonchaos transition as $W_{\gamma}$ increases. In order to quantitatively depict the localization of the Krylov wave function, we employ the inverse participation ratio (IPR), defined as
$I_{K}=\sum_{n}\left|\varphi_{n}(t)\right|^{4}.$ (4)
The decreasing of $I_{K}$ indicates delocalization of $\varphi_{n}$, and its values at $t=300$ and $t=1000$ are shown in Figs. 2(d) and 2(e).
We identify three distinct regimes for the dissipation strength $W_{\gamma}$, separated by two critical values $W_{L}$ and $W_{C}$. The salient features of the complexity growth $C_{K}(t)$ and the localization of the Krylov wave function, characterized by the IPR, are contrasted in Table 1. Universally across all regimes, the short-time growth of complexity is quadratic, $C_{K}(t)\sim(1+W_{\gamma}^{2})t^{2}$ [dashed line in Fig. 2(a)]; fitting details are provided in Appendix C. The absence of an exponential growth indicates a departure from the operator growth hypothesis. Combined with the nonlinear Lanczos coefficients shown in Fig. 3, this precludes a chaos diagnosis from early-time dynamics in our model. We therefore focus on the intermediate- and long-time behavior.
In the weak dissipative regime ($W_{\gamma}<W_{L}$), $C_{K}$ develops a pronounced plateau at intermediate times, whose height rises and duration shrinks as $W_{\gamma}$ increases [red curves in Fig. 2(a)]. This signifies a prethermal stage where the Krylov wave function $\varphi_{n}$ is tightly confined at early times [Fig. 1(d), left], leading to an extraordinarily high IPR at $t=300$ [Fig. 2(c)] that decreases only weakly with system size. Chaotic dynamics are not firmly established due to this mid-term plateau, but are expected to emerge in the long-time regime after $C_{K}$ abruptly increases to a high value, which is nearly independent of $W_{\gamma}$. Concurrently, $I_{K}$ decreases as the Krylov wave function finally spreads across the entire Krylov space. Thus, diagnosing chaos in the weak dissipative regime requires observing the long-time evolution of the complexity.
For moderate dissipation ($W_{L}<W_{\gamma}<W_{C}$), the prethermal plateau disappears. The intermediate growth of $C_{K}$ accelerates with increasing $W_{\gamma}$, yet its ultimate saturation value diminishes [green to yellow lines in Fig. 2(a)]. This behavior forms a linear envelope, indicated by the dotted line in the same figure. In contrast to Ref. [26], where a linear ramp in complexity is linked to spectral rigidity and signifies quantum chaos, our results reflect chaos through a different form of complexity growth. This difference is also mirrored in the Lanczos coefficients. A linear ramp of $C_{K}$ results from $b_{n}\propto\sqrt{1-n/2^{L}}$, a special case of Tsallis $q$-log statistics with $q=0$ [46, 41]. Our model, however, yields $q>0$, as shown in Fig. 3 and Appendix D. We also note that the suppression of the saturated complexity, the enhanced IPR, and the confinement of the Krylov wave function to $n<O(L^{2})$ [Fig. 1(d), middle] suggest a weak breaking of ergodicity. This makes the chaotic dynamics for $W_{L}<W_{\gamma}<W_{C}$ distinct from those in the weak dissipation regime. The saturated $C_{K}$ exhibits a transitional scaling with system size [Fig. 2(c)], the timescale to reach saturation is much shorter, and the chaotic evolution is restricted to a smaller Krylov subspace.
In the strongly dissipative regime ($W_{\gamma}>W_{C}$), $C_{K}$ saturates rapidly to a low value [blue lines in Fig. 2(a)]. The Krylov wave function remains localized to $n<O(L)$ [Fig. 1(d), right], and the IPR is higher than that in the other regimes. These features signify a marked departure from conventional quantum chaos. As we will show later, the complex level statistics also confirm that $W_{C}$ marks the transition from chaotic to nonchaotic phases.
Figure 3: Lanczos coefficients for $L=10$ and $W_{\gamma}=$ (a) 0.0118, (b) 0.2, (c) 1.0, and (d) 3.0. Upper panels show $\langle|b_{n}|\rangle$ (solid blue) and $\langle|c_{n}|\rangle$ (dashed red), with dotted lines representing $\sim n$ for $n<L$, and $\sim\sqrt{1-n/2^{L}}$ for $n>L$. Lower panels show $\langle\cos\theta_{n}\rangle$.
IV Krylov variance and reciprocity